The production cost, represented by y, of smart phone devices are equal to the product of computing components a and the square of the number of devices x, added to the product pf labour b and number of devices x; after which savings on reusable material c are subtracted.


2.1. Write down an equation to represent the production cost y in terms of a, x, b and c


2.2. Change the subject of the equation/formula in 2.1 to:


2.2.1.a=


b=


c=


2.2.2 Determine the production cost of 7 devices, if labour amounts to R4140, computing components cost is R1035 and savings on reusable material is R1725

1
Expert's answer
2022-02-23T12:28:45-0500

2.1) "y=ax^2+bx-c"

2.2) "a= \frac{y-bx+c}{x^2}"


"b=\frac{ y-ax^2+c}{x}"


"c=ax^2+bx-y"


"2.2.2)\>y=1035(7^2)+4140(7)-1725\\=R\>77970"